OpenStudy (gabylovesyou):

What is the equation of the line that passes through the point (1,6) and has a slope of 3?

OpenStudy (gabylovesyou):

@jim_thompson5910

OpenStudy (anonymous):

The equation has the form: y=ax+b a is 3 in this case, so y=3x+b Now just put the x and y from that point into the equation and b must be a number that satisfy it. 6=3*1+b

jimthompson5910 (jim_thompson5910):

y = mx+b 6 = (3)(1)+b 6 = 3+b

jimthompson5910 (jim_thompson5910):

Keep going to solve for b

OpenStudy (gabylovesyou):

b = 3

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

So the equation is y = 3x + 3

OpenStudy (gabylovesyou):

so is the equation y = 3x + 3?

jimthompson5910 (jim_thompson5910):

yes you nailed it

OpenStudy (gabylovesyou):

lol yea